OpenStudy (anonymous):

whic expression is equal to (fg)(x) f(x)=x+6; g(x)=x^2-2 a.x^3-x^2-6x-12 b.x^3+6x^2-2x-12 c. x^3-12 d. x^3+4

OpenStudy (pawanyadav):

For fg(x) Put the function g(x) in place of x in function f(x) then you will get fg(x) Example: f(x)=x+1, g(x)=x^2 Then for fg(x) put g(x) or x^2 in place of x So. fg(x)=x^2+1